### 06 Mar 2026 v1

@@ 1 Introduction
From a probabilistic perspective, DA can be formulated as a Bayesian filtering problem [30], where the goal is to sequentially infer the posterior distribution of the system state conditioned on all available observations. Among approximate Bayesian filtering methods, the ensemble Kalman filter (EnKF) [17] has become one of the most widely adopted techniques due to its favorable computational scaling, derivative-free implementation, and ensemble-based representation of uncertainty. For linear systems with Gaussian noise, the EnKF provides a Monte Carlo approximation of the classical Kalman filter and converges to the exact Bayesian solution as the ensemble size increases. However, in realistic applications involving strongly nonlinear dynamics, partial observations, and non-Gaussian error statistics, the EnKF often becomes suboptimal and may even suffer from filter divergence [20]. Importantly, this degradation is not intrinsic to the Kalman update itself, but rather stems from a structural mismatch between the linear–Gaussian assumptions underlying the EnKF analysis step and the true nonlinear dynamics and observation processes. As a result, the posterior ensemble is effectively constrained to an affine subspace determined by local linearizations of the model and observation operator [30, 27], leading to biased estimates and loss of stability.
Extensive research has aimed to enhance the robustness of EnKF-type methods in nonlinear and non-Gaussian settings. Sigma-point filters, including the unscented and cubature Kalman filters, propagate carefully chosen samples through nonlinear models to improve moment approximations [18, 3]. Particle filters offer greater flexibility by representing the posterior with weighted ensembles [30], but they typically suffer from severe weight degeneracy in high-dimensional systems. Alternative approaches reinterpret the Bayesian update as a deterministic transport or flow of probability measures [31], while hybrid variational–ensemble methods incorporate gradient information to better handle nonlinear observation operators [9]. Despite these advances, achieving accurate, stable, and scalable filtering for strongly nonlinear, high-dimensional systems remains a fundamental challenge.
@@
To address these issues, Chen et al. [15] proposed the multi-domain encoder–decoder latent assimilation (MEDLA) framework, which maps heterogeneous observations into a unified latent space using multiple encoders and decoders. While MEDLA improves flexibility under diverse observation settings, its latent dynamics are governed by unconstrained nonlinear neural operators, which can limit interpretability and long-term stability. In parallel, Koopman operator theory offers a complementary perspective by representing nonlinear dynamical systems through linear evolution in a suitably lifted feature space [8, 26]. Koopman autoencoders operationalize this idea by learning encoder–decoder pairs such that latent variables evolve under a learned linear operator [5, 28]. Crucially, linear latent dynamics provide improved interpretability, numerical stability, and structural compatibility with Kalman filtering.
@@
On the theoretical side, we analyze the proposed framework under a smooth manifold assumption and establish generalization error bounds for learning stable linear latent dynamics. Extensive numerical experiments on representative nonlinear and chaotic systems show that the proposed LAE-EnKF consistently improves assimilation accuracy, stability, and robustness compared with standard EnKF and existing autoencoder-assisted approaches, while maintaining favorable computational efficiency.
@@ 3.1 Latent autoencoder assimilation
where $\mathbf{A}\in\mathbb{R}^{n\times n}$ is a trainable matrix. This construction is closely related to Koopman-based representations [5, 28], in which nonlinear dynamics are lifted to linear evolution in a learned feature space. Crucially, the explicit linear structure of (5) is not merely a modeling convenience, but a deliberate design choice that enables structural compatibility with Kalman filtering.
@@
Here $\mathbf{H}\in\mathbb{R}^{m\times n}$ is a prescribed or learned observation matrix. By embedding both states and observations into a common latent space, we obtain a closed linear state-space system in the latent variables. This construction eliminates the latent-space mismatch that arises when state evolution and observation mapping are learned independently, and provides a consistent foundation for Kalman-type updates.
@@ 3.2 Network architecture and learning framework
The latent autoencoder (LAE) framework consists of a state encoder $\mathscr{E}$, a decoder $\mathscr{D}$, a latent linear transition operator $\mathbf{A}$, and an observation encoder $\mathscr{E}_{\mathrm{obs}}$, as illustrated in Fig. 1. The central objective is to learn a low-dimensional representation that is expressive enough to capture nonlinear system dynamics, while enforcing a structure that is compatible with linear–Gaussian filtering. In particular, the learned latent representation is designed so that the system evolution and observation process admit linear models, thereby restoring structural consistency with ensemble Kalman filtering.
@@
$+\lambda_{\mathrm{latent}}\|\mathbf{A}\mathscr{E}(\mathbf{x}_{k}^{(i)})-\mathscr{E}(\mathbf{x}_{k+1}^{(i)})\|_{2}^{2}\Big)$
(8) $+\lambda_{\mathrm{reg}}R(\mathbf{A}),$
@@
$R(\mathbf{A})=\bigl(\max\{0,\|\mathbf{A}\|_{2}-1\}\bigr)^{2}$
@@
[…]
Figure 1: The architecture of latent autoencoder framework. It consists of four principal trainable components: a state encoder $\mathscr{E}$ that projects physical states into a compact latent space, a decoder $\mathscr{D}$ that reconstructs physical states from latent representations, a linear transition operator $\mathbf{A}$ that governs latent dynamics, and an observation encoder $\mathscr{E}_{\text{obs}}$ that maps observations into the same latent manifold.
@@
After the latent dynamics have been identified, the encoder $\mathscr{E}$, decoder $\mathscr{D}$, and transition operator $\mathbf{A}$ are kept fixed. In the second stage, an observation encoder $\mathscr{E}_{\mathrm{obs}}$ is trained to embed observations into the same latent coordinate system. We consider a linear observation model in the latent space and define the observation alignment loss
(9) $\mathcal{L}_{\mathrm{obs}}(\mathscr{E}_{\mathrm{obs}})=\frac{1}{N_{\mathrm{traj}}(K+1)}\sum_{i=1}^{N_{\mathrm{traj}}}\sum_{k=0}^{K}\|\mathscr{E}_{\mathrm{obs}}(\mathbf{y}_{k}^{(i)})-\mathbf{H}\mathscr{E}(\mathbf{x}_{k}^{(i)})\|_{2}^{2},$
where $\mathbf{H}$ denotes a prescribed linear observation operator in the latent space. In practice, a convenient and commonly used choice is $\mathbf{H}=\mathbf{I}$, in which case the observation encoder directly maps measurements into the latent state space without introducing additional trainable parameters.
@@ 3.3 Latent autoencoder ensemble Kalman filter
10:  Compute latent observation: $\widehat{\mathbf{y}}_{k}^{(j)}=\mathbf{H}\mathbf{z}_{k|k-1}^{(j)},\quad j=1,\cdots,N_{e}$.
@@ 4 Theoretical Analysis
Let $\mathcal{M}\subset\mathbb{R}^{D}$ be a compact $n$-dimensional $C^{2}$ submanifold. Assume that $\mathcal{M}$ admits a global bi-Lipschitz parameterization $\varphi:\mathcal{M}\rightarrow[-\Lambda,\Lambda]^{n}$ with inverse $\phi=\varphi^{-1}:\varphi(\mathcal{M})\rightarrow\mathcal{M}$. That is, there exist constants $C_{\varphi},C_{\phi}>0$ such that for all $\mathbf{v}_{1},\mathbf{v}_{2}\in\mathcal{M}$,
@@
Let $\{(\mathbf{x},\mathbf{x}^{+})\}_{k=1}^{N}$ denote the input-target pairs, where $\mathbf{x}_{k}^{+}=\mathbf{F}(\mathbf{x}_{k})$ represents the one-step evolution under the true dynamics. We assume that these pairs are sampled i.i.d. according to the stationary transition distribution:
(10) $\left(\mathbf{x},\mathbf{x}^{+}\right)\overset{\text{i.i.d.}}{\sim}\mathbb{P}_{X,X^{+}},$
supported on $\mathcal{M}\times\mathcal{M}$.
For the loss function (3.2), the weights $\lambda_{\mathrm{rec}},\lambda_{\mathrm{pred}},\lambda_{\mathrm{latent}},\lambda_{\mathrm{reg}}$ are tuning parameters used in practice to balance the relative scales of reconstruction, one-step prediction, latent consistency, and stability regularization. For the theoretical analysis, these weights act only as positive multiplicative scalings of the loss terms and therefore affect the generalization bound only through the leading constant, without changing the convergence rate. Accordingly, in the theoretical part, we set $\lambda_{\mathrm{rec}}=\lambda_{\mathrm{pred}}=\lambda_{\mathrm{latent}}=\lambda_{\mathrm{reg}}=1$ and use the following simplified loss function:
$\mathcal{L}_{\mathrm{lat}}^{\mathrm{pop}}(\mathscr{E},\mathscr{D},\mathbf{A}):=\mathbb{E}_{(\mathbf{x},\mathbf{x}^{+})\sim\mathbb{P}_{X,X^{+}}}\Big[$ $\|\mathscr{D}(\mathbf{A}\mathscr{E}(\mathbf{x}))-\mathbf{F}(\mathbf{x})\|_{2}^{2}$
$+\|\mathscr{D}(\mathscr{E}(\mathbf{x}))-\mathbf{x}\|_{2}^{2}$
(11) $+\|\mathbf{A}\mathscr{E}(\mathbf{x})-\mathscr{E}(\mathbf{F}(\mathbf{x}))\|_{2}^{2}+R(\mathbf{A})\Big].$
Under these assumptions, we obtain the following finite-sample upper bound for the squared generalization error of the LAE.
@@
(12) $\mathbb{E}_{\mathcal{S}}\!\left[\mathcal{L}_{\mathrm{lat}}^{\mathrm{pop}}(\widehat{\mathscr{E}},\widehat{\mathscr{D}},\widehat{\mathbf{A}})\right]\leq CD^{2}\log^{2}DN^{-\frac{2}{n+2}}\log^{4}N,$
for some constant $C>0$ depending only on $n$, $B$, $\Lambda$, $\rho$, the Lipschitz constants of the parametrization maps, and the volume of $\mathcal{M}$.
A detailed proof is provided in Appendix A. Our analysis follows the general proof strategy developed for the single-chart autoencoder framework [25], where the three components of the population loss (4) are bounded separately. Compared with [25], our setting includes an additional latent linear evolution term $\mathbf{A}$, which introduces an extra intermediate factor of order $\mathcal{O}(N^{-\frac{2}{n+2}})$ in the estimation error. However, this additional term does not affect the overall convergence rate. Consequently, the final squared generalization error remains $\mathcal{O}(N^{-\frac{2}{n+2}}\log^{4}N)$ matching the rate achieved in the single-chart autoencoder setting.
The following theorem provides a corresponding generalization bound for the observation encoder under similar architectural assumptions. For simplicity of exposition, we consider the latent observation operator is fixed to the identity, i.e., $\mathbf{H}=\mathbf{I}$.
@@
Suppose that Assumption 1 holds, and that the pre-trained components $\mathscr{E}$, $\mathscr{D}$, and $\mathbf{A}$ are fixed. Let $\widehat{\mathscr{E}}_{\mathrm{obs}}$ denote a global minimizer of the observation alignment objective (9). Assume that the observation encoder class has sufficient approximation capacity to represent the mapping $\mathbf{y}\mapsto\mathscr{E}(\mathbf{x})$ of the data distribution induced by the observation process. Then the expected generalization error satisfies
(13) $\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(\mathbf{x},\mathbf{y})}\!\left[\bigl\|\widehat{\mathscr{E}}_{\mathrm{obs}}(\mathbf{y})-\mathscr{E}(\mathbf{x})\bigr\|_{2}^{2}\right]\leq C\,D_{y}^{2}\log^{2}D_{y}\,N^{-\frac{2}{n+2}}\log^{4}N,$
for some constant $C>0$ depending only on $n$, $B$, $\Lambda$, and volume of the manifold $\mathcal{M}$.
The proof follows similar arguments as in Theorem 4.1 and is therefore omitted for brevity.

### 29 Apr 2026 v2

@@ 1 Introduction
From a probabilistic perspective, DA can be formulated as a Bayesian filtering problem [37], where the goal is to sequentially infer the posterior distribution of the system state conditioned on all available observations. Among approximate Bayesian filtering methods, the ensemble Kalman filter (EnKF) [22] has become one of the most widely adopted techniques due to its favorable computational scaling, derivative-free implementation, and ensemble-based representation of uncertainty. In the linear–Gaussian setting, the EnKF recovers the classical Kalman filter, which provides the optimal linear update for Gaussian systems, in the large-ensemble limit [34]. For nonlinear systems, however, the large-ensemble limit of the EnKF generally does not recover the true filtering distribution [21]. This limitation has been investigated from various theoretical perspectives, including mean-field and non-asymptotic analyses, which provide insight into the behavior of ensemble-based approximations in nonlinear settings [14, 2, 3, 39]. Despite these theoretical insights, the practical performance of the EnKF remains fundamentally constrained by finite ensemble sizes. Sampling errors in the empirical covariance can induce spurious long-range correlations, significantly degrading estimation accuracy. To alleviate these issues, techniques such as localization and covariance inflation are commonly employed [43, 44]. However, these remedies address sampling errors but not the underlying limitation. In nonlinear, partially observed, or non-Gaussian settings, the EnKF can become suboptimal and even diverge [25]. This degradation is not intrinsic to the Kalman update itself, but rather stems from a structural mismatch between the linear-Gaussian assumptions underlying the EnKF analysis step and the true nonlinear dynamics and observation processes. As a result, the posterior ensemble is effectively constrained to an affine subspace determined by local linearizations of the model and observation operator [37, 33], leading to biased estimates and loss of stability.
Extensive research has aimed to enhance the robustness of EnKF-type methods in nonlinear and non-Gaussian settings. Sigma-point filters, including the unscented and cubature Kalman filters, propagate carefully chosen samples through nonlinear models to improve moment approximations [23, 5]. Particle filters offer greater flexibility by representing the posterior with weighted ensembles [37], but they typically suffer from severe weight degeneracy in high-dimensional systems. More recently, generative approaches based on score and diffusion models have been proposed to directly model complex posterior distributions [8, 30, 38]. Alternative approaches reinterpret the Bayesian update as a deterministic transport or flow of probability measures [40], while hybrid variational–ensemble methods incorporate gradient information to better handle nonlinear observation operators [12]. Despite these advances, achieving accurate, stable, and scalable filtering for strongly nonlinear, high-dimensional systems remains a fundamental challenge.
@@
To address these issues, Chen et al. [19] proposed the multi-domain encoder-decoder latent assimilation (MEDLA) framework, which maps heterogeneous observations into a unified latent space using multiple encoders and decoders. While MEDLA improves flexibility under diverse observation settings, its latent dynamics are governed by unconstrained nonlinear neural operators, which can limit interpretability and long-term stability. In parallel, Koopman operator theory offers a complementary perspective by representing nonlinear dynamical systems through linear evolution in a suitably lifted feature space [11, 32]. Koopman autoencoders operationalize this idea by learning encoder–decoder pairs such that latent variables evolve under a learned linear operator [7, 35]. Crucially, linear latent dynamics provide improved interpretability, numerical stability, and structural compatibility with Kalman filtering.
@@
On the theoretical side, we analyze the proposed framework under a smooth manifold assumption and establish generalization error bounds for both the latent linear dynamics and the observation encoder, while explicitly accounting for the effects of dynamical perturbations and observation noise. Extensive numerical experiments on representative nonlinear and chaotic systems show that the proposed LAE-EnKF consistently improves assimilation accuracy, stability, and robustness compared with standard EnKF and existing autoencoder-assisted approaches, while maintaining favorable computational efficiency.
@@ 2 Data Assimilation
In this section, we introduce the problem formulation and the EnKF framework. For ease of presentation, we adopt the following notation. Lower-case letters denote scalars, bold lower-case letters denote vectors, and upper-case letters denote matrices. Calligraphic letters represent manifolds, sets, and function classes, while script letters denote neural network mappings. We use $\widehat{\cdot}$ to denote empirical or learned estimators, and $\widetilde{\cdot}$ to denote latent variables or approximation functions arising in the theoretical analysis.
@@ 3.1 Latent autoencoder assimilation
where $\mathbf{A}\in\mathbb{R}^{n\times n}$ is a trainable linear operator, parameterized as a weight matrix and learned jointly with the encoder and decoder through gradient-based optimization. This construction is closely related to Koopman-based representations [7, 35], in which nonlinear dynamics are lifted to linear evolution in a learned feature space. Crucially, the explicit linear structure of (5) is not merely a modeling convenience, but a deliberate design choice that enables structural compatibility with Kalman filtering.
@@
Here $\mathbf{H}\in\mathbb{R}^{m\times n}$ denotes a linear observation operator in the latent space, which can either be specified a priori or learned jointly with the observation encoder $\mathscr{E}_{\mathrm{obs}}$. By embedding both states and observations into a common latent space, we obtain a closed linear state-space system in the latent variables. This construction eliminates the latent-space mismatch that arises when state evolution and observation mapping are learned independently, and provides a consistent foundation for Kalman-type updates.
@@ 3.2 Network architecture and learning framework
The latent autoencoder (LAE) framework consists of a state encoder $\mathscr{E}$, a decoder $\mathscr{D}$, a latent linear transition operator $\mathbf{A}$, an observation encoder $\mathscr{E}_{\mathrm{obs}}$, and a latent observation operator $\mathbf{H}$, as illustrated in Fig. 1. The central objective is to learn a low-dimensional representation that is expressive enough to capture nonlinear system dynamics, while enforcing a structure that is compatible with linear–Gaussian filtering. In particular, the learned latent representation is designed so that the system evolution and observation process admit linear models, thereby restoring structural consistency with ensemble Kalman filtering.
[…]
Figure 1: The architecture of latent autoencoder framework. It consists of a state encoder $\mathscr{E}$ that maps physical states to a latent space, a decoder $\mathscr{D}$ that reconstructs states from latent representations, a linear transition operator $\mathbf{A}$ that governs latent dynamics, an observation encoder $\mathscr{E}_{\mathrm{obs}}$ that embeds observations into the latent space, and a latent observation operator $\mathbf{H}$, which may be specified a priori or learned.
@@
(8) $+\lambda_{\mathrm{latent}}\|\mathbf{A}\mathscr{E}(\mathbf{x}_{k}^{(i)})-\mathscr{E}(\mathbf{x}_{k+1}^{(i)})\|_{2}^{2}\Big)+\lambda_{\mathrm{reg}}R(\mathbf{A}),$
@@
$R(\mathbf{A})=\bigl(\max\{0,\|\mathbf{A}\|_{2}-1\}\bigr)^{2},$
@@
After the latent dynamics have been identified, the encoder $\mathscr{E}$, decoder $\mathscr{D}$, and transition operator $\mathbf{A}$ are kept fixed. In the second stage, an observation encoder $\mathscr{E}_{\mathrm{obs}}$ is trained to embed observations into the same latent coordinate system. We consider a linear observation model in the latent space with operator $\mathbf{H}$, which can either be specified a priori or learned jointly with $\mathscr{E}_{\mathrm{obs}}$, and define the observation alignment loss
(9) $\mathcal{L}_{\mathrm{obs}}(\mathscr{E}_{\mathrm{obs}})=\frac{1}{N_{\mathrm{traj}}(K+1)}\sum_{i=1}^{N_{\mathrm{traj}}}\sum_{k=0}^{K}\|\mathscr{E}_{\mathrm{obs}}(\mathbf{y}_{k}^{(i)})-\mathbf{H}\mathscr{E}(\mathbf{x}_{k}^{(i)})\|_{2}^{2}\,.$
This objective enforces consistency between the observation encoder and the latent representation induced by $\mathscr{E}$ and $\mathbf{H}$. It can be interpreted as a regression objective toward the latent observation model, rather than a reconstruction requirement. In particular, it does not require observations to uniquely determine the underlying state, but only to produce a latent embedding compatible with the linear observation structure. When $\mathbf{H}$ is learned jointly with $\mathscr{E}_{\mathrm{obs}}$, a norm-based regularization on $\mathbf{H}$ is needed to mitigate scaling ambiguity and to ensure the stability.
@@ 3.3 Latent autoencoder ensemble Kalman filter
10:  Compute latent observation: $\widehat{\mathbf{y}}_{k}^{(j)}=\mathbf{H}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\widehat{\mathbf{z}}}_{k|k-1}^{(j)},\quad j=1,\cdots,N_{e}$.
@@ 4 Theoretical Analysis
Let $\mathcal{M}\subset\mathbb{R}^{D}$ be a compact $n$-dimensional $C^{2}$ submanifold. Assume that $\mathcal{M}$ admits a global bi-Lipschitz parameterization $\varphi:\mathcal{M}\rightarrow[-\Lambda,\Lambda]^{n}$ for some $\Lambda>0$, with inverse $\phi=\varphi^{-1}:\varphi(\mathcal{M})\rightarrow\mathcal{M}$. That is, there exist constants $C_{\varphi},C_{\phi}>0$ such that for all $\mathbf{v}_{1},\mathbf{v}_{2}\in\mathcal{M}$,
@@
Let $\{(\mathbf{x},\mathbf{x}^{+})\}_{k=1}^{N}$ denote the input-target pairs, where $\mathbf{x}_{k}^{+}=\mathbf{F}(\mathbf{x}_{k})+\mathbf{w}_{k}$ and $\mathbf{w}_{k}$ denotes an additive perturbation term, assumed to be i.i.d., independent of $\mathbf{x}_{k}$, with $\mathbb{E}[\mathbf{w}_{k}]=0$ and $\mathbb{E}\|\mathbf{w}_{k}\|_{2}^{2}<\infty$. We assume that $(\mathbf{x},\mathbf{x}^{+})$ are sampled i.i.d. from a joint distribution $\mathbb{P}_{X,X^{+}}$ supported on a neighborhood of $\mathcal{M}\times\mathcal{M}$.
For the loss function (3.2), the weights $\lambda_{\mathrm{rec}},\lambda_{\mathrm{pred}},\lambda_{\mathrm{latent}},\lambda_{\mathrm{reg}}$ are tuning parameters used in practice to balance the relative scales of reconstruction, one-step prediction, latent consistency, and stability regularization. From a theoretical perspective, they act as positive multiplicative constants and can be absorbed into the leading constants of the generalization bound, and hence do not affect the convergence rate. Without loss of generality, we set $\lambda_{\mathrm{rec}}=\lambda_{\mathrm{pred}}=\lambda_{\mathrm{latent}}=\lambda_{\mathrm{reg}}=1$ and consider the simplified population loss:
$\mathcal{L}_{\mathrm{lat}}^{\mathrm{pop}}(\mathscr{E},\mathscr{D},\mathbf{A}):=\mathbb{E}_{(\mathbf{x},\mathbf{x}^{+})\sim\mathbb{P}_{X,X^{+}}}\Big[$ $\|\mathscr{D}(\mathbf{A}\mathscr{E}(\mathbf{x}))-\mathbf{F}(\mathbf{x}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\mathbf{w}}\|_{2}^{2}+\|\mathscr{D}(\mathscr{E}(\mathbf{x}))-\mathbf{x}\|_{2}^{2}$
(10) $+\|\mathbf{A}\mathscr{E}(\mathbf{x})-\mathscr{E}(\mathbf{F}(\mathbf{x}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\mathbf{w}})\|_{2}^{2}+R(\mathbf{A})\Big].$
Let $\mathcal{S}=\{(\mathbf{x}_{i},\mathbf{x}_{i}^{+},\mathbf{y}_{i})\}_{i=1}^{N}$ denote i.i.d. samples drawn from a joint distribution $\mathbb{P}_{X,X^{+},Y}$ induced by the underlying dynamics and the observation model. Under these assumptions, we obtain the following finite-sample upper bound for the squared generalization error of the LAE.
@@
(11) $\mathbb{E}_{\mathcal{S}}\!\left[\mathcal{L}_{\mathrm{lat}}^{\mathrm{pop}}(\widehat{\mathscr{E}},\widehat{\mathscr{D}},\widehat{\mathbf{A}})\right]\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{0}}D^{2}\log^{2}DN^{-\frac{2}{n+2}}\log^{4}N+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}(1+C_{1}^{2}C^{2}_{\varphi})\mathbb{E}\|\mathbf{w}\|_{2}^{2}},$
for some constant ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{0}}>0$ depending only on $n$, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}$}, $B$, $\Lambda$, $\rho$, the Lipschitz constants of the parametrization maps, and the volume of $\mathcal{M}$.
A detailed proof is provided in Appendix A. The upper bound contains the term $\mathbb{E}\|\mathbf{w}\|_{2}^{2}$, which represents the irreducible error induced by stochastic perturbations in the dynamics. For deterministic systems, where the evolution is noise-free, this term vanishes. Our analysis follows the general proof strategy developed for the single-chart autoencoder framework [31], where the three components of the population loss (4) are bounded separately. Compared with [31], our setting includes an additional latent linear evolution term $\mathbf{A}$, which introduces an extra intermediate factor of order $\mathcal{O}(N^{-\frac{2}{n+2}})$ in the estimation error. However, this additional term does not affect the overall convergence rate. Consequently, the squared generalization error consists of an estimation error term of order $\mathcal{O}\bigl(N^{-\frac{2}{n+2}}\log^{4}N\bigr)$ together with an irreducible error term due to the stochastic perturbations in the dynamics.
The following theorem establishes a generalization bound for the observation encoder $\mathscr{E}_{obs}$ defined above, which aims to approximate the latent representation $\mathbf{H}\,\mathscr{E}(\mathbf{x})$ from observations $\mathbf{y}$. To ensure identifiability and stability of this mapping, we assume that the observation operator $\mathcal{H}$ satisfies a local stability condition.
@@
Suppose that Assumption 1 holds, and that the pre-trained components $\mathscr{E}$, $\mathscr{D}$, and $\mathbf{A}$ are fixed. Let $\widehat{\mathscr{E}}_{\mathrm{obs}}$ denote a global minimizer of the observation alignment objective (9). Assume that the observation encoder class has sufficient approximation capacity to represent the regression function $\mathbf{y}\;\mapsto\;\mathbb{E}[\mathbf{H}\mathscr{E}(\mathbf{x})\mid\mathbf{y}]$, under the data distribution induced by the observation process. Then the expected generalization error satisfies
(12) $\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(\mathbf{x},\mathbf{y})}\left[\bigl\|\widehat{\mathscr{E}}_{\mathrm{obs}}(\mathbf{y})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}}\mathscr{E}(\mathbf{x})\bigr\|_{2}^{2}\right]\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{0}^{obs}}D_{y}^{2}\log^{2}D_{y}N^{-\frac{2}{n+2}}\log^{4}N{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+C_{1}^{obs}\mathbb{E}\|\boldsymbol{\eta}\|_{2}^{2}},$
where the constant $C_{0}^{obs}>0$ depends on $n$, $R_{obs}$, $\Lambda$, and the volume of $\mathcal{M}$, and $C_{1}^{obs}>0$ depends on $\|\mathbf{H}\|_{2}$, $C_{\phi}$ and the stability constant $C_{\mathcal{H}}$.
A detailed proof is provided in Appendix B. The error bound consists of an estimation term of order $\mathcal{O}\bigl(N^{-\frac{2}{n+2}}\log^{4}N\bigr)$ and an irreducible term induced by the observation noise $\boldsymbol{\eta}$. The rate is governed by the intrinsic dimension $n$ and matches that in Theorem 4.1.